Find energy of each of the photons which (i) correspond to light of frequency 3×1015 Hz. (ii) have wavelength of 0.50 Å.

Grade:12

1 Answers

Gaurav
9 years ago
(i) Energy (E) of a photon is given by the expression, E = hv
Where, h = Planck’s constant = 6.626 × 10–34 Js
ν = frequency of light = 3 × 1015 Hz
Substituting the values in the given expression of E:
E = (6.626 × 10–34) (3 × 1015) E = 1.988 × 10–18 J

(ii) Energy (E) of a photon having wavelength (λ) is given by the expression, h = Planck’s constant = 6.626 × 10–34 Js
c = velocity of light in vacuum = 3 × 108 m/s Substituting the values in the given expression of E:
E = 3.98 X 1015 J